OpenStudy (anonymous):

7x - 3y = 4 2x - 4y = 1 Which of the following systems is not the same as solving the system shown? 28x - 12y = 16 and -6x + 12y = -3 14x - 6y = 4 and -14x + 28y = 1 -28x + 12y = -16 and 28x - 56y = 14

OpenStudy (anonymous):

do u wanna an explanation or the final answer directly ?

OpenStudy (anonymous):

the last Answer (-28x+12y=16 & 28x-56y=14 )

OpenStudy (anonymous):

could you please help? 3x - 4y = 7 3x + 2y = -5 When the second equation is subtracted from the first, the resulting equation is _____.